OpenStudy (anonymous):

A= (-6, -3) C= (-3, -5) If point C is the midpoint of AB, what is the coordinate of point B? (-3, 1) (-9, 2) (-1, -8) (0, -7)

OpenStudy (amistre64):

C+C-A

hero (hero):

(0,-7)

OpenStudy (amistre64):

C= (-3, -5) +C= (-3, -5) -A= ( 6, 3) ----------- 0,-7

hero (hero):

amistre, wth?

hero (hero):

What kind of method is that?

OpenStudy (anonymous):

Thank ya'll so much! I couldn't figure it out

hero (hero):

Explain in more detail please

OpenStudy (amistre64):

well, C-A is distance and direction to midpoint, added to the midpoint C is your end point

OpenStudy (amistre64):

hence: C+(C-A) = other end

hero (hero):

Is that the best explanation you have?

OpenStudy (amistre64):

yep. distance from A to C added to C

hero (hero):

You're not good with the whole "general explanation" thing are you?

OpenStudy (amistre64):

not really, i need a better thesaurus :)

hero (hero):

amistre, you come up with the strangest mathematical equations, but can't explain to save your life? Are you the reincarnation of tesla, einstein, and newton?

OpenStudy (amistre64):

and i guy named Fred from the 300BCs

hero (hero):

i guy?

OpenStudy (amistre64):

he cant type lol

hero (hero):

Funny